OpenStudy (anonymous):

the length of a rectangle is 6 meters more than its width. if the area is 91 square meters, what is the length and width?

OpenStudy (anonymous):

this is in regards to quadratic equatons

OpenStudy (anonymous):

l=w+6 w*l=91 w(w+6)=91 w^2+6w-91=0 now solve...

OpenStudy (anonymous):

ok thank u!:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I cant find a common factor for 91 and 6?

OpenStudy (anonymous):

got it:)

OpenStudy (campbell_st):

if you can't find factors use the general quadratic formula... less stress..

OpenStudy (campbell_st):

but I think its 13 and 7

OpenStudy (campbell_st):

oops 13 and - 7

OpenStudy (anonymous):

yes thank you:) I got (w-7)(w+13)

OpenStudy (campbell_st):

don't look for common factors.. look for factors of the constant that add to the coefficient of x

OpenStudy (anonymous):

would this be square meters as the units?

OpenStudy (anonymous):

take only the positive value,the unit is meters

OpenStudy (anonymous):

ok thank you:) so, look for factors of the constant that add to the coeffiecent of x :)..gotcha:)

OpenStudy (anonymous):

ok thx:)

OpenStudy (anonymous):

w ( w +6) -91 = 0 w^2 + 6w - 91 = 0 x = -3 +10 = 7 (omit negative value)